# SFGNET: SEMANTIC AND FREQUENCY GUIDED NETWORK FOR CAMOUFLAGED OBJECT DETECTION

### v1

@@ Abstract
Camouflaged object detection (COD) aims to segment objects that blend into their surroundings. However, most existing studies overlook the semantic differences among textual prompts of different targets as well as fine-grained frequency features. In this work, we propose a novel Semantic and Frequency Guided Network (SFGNet), which incorporates semantic prompts and frequency-domain features to capture camouflaged objects and improve boundary perception. We further design Multi-Band Fourier Module(MBFM) to enhance the ability of the network in handling complex backgrounds and blurred boundaries. In addition, we design an Interactive Structure Enhancement Block (ISEB) to ensure structural integrity and boundary details in the predictions. Extensive experiments conducted on three COD benchmark datasets demonstrate that our method significantly outperforms state-of-the-art approaches. Our code is publicly available at https://github.com/winter794444/SFGNetICASSP2026.
@@ 3.1.1 Datasets
PFNet None ResNet 0.782 0.695 0.085 0.855 0.800 0.660 0.040 0.877 0.829 0.745 0.053 0.887
ZoomNet None ResNet 0.820 0.752 0.066 0.877 0.838 0.729 0.029 0.888 0.853 0.784 0.043 0.896
FSPNet None ViT 0.851 0.802 0.056 0.905 0.850 0.755 0.028 0.912 0.879 0.816 0.035 0.914
FEDER None Res2Net 0.836 0.807 0.066 0.897 0.844 0.748 0.029 0.911 0.862 0.824 0.042 0.913
HitNet None PVT 0.849 0.809 0.055 0.906 0.871 0.806 0.023 0.935 0.875 0.834 0.037 0.926
FSEL None PVT 0.885 0.857 0.040 0.942 0.877 0.799 0.021 0.937 0.892 0.852 0.030 0.941
CamoFormer None PVT 0.872 0.831 0.046 0.929 0.869 0.786 0.023 0.932 0.892 0.847 0.030 0.939
FHCNet None Swin+Res2Net 0.885 0.842 0.039 0.934 0.869 0.778 0.023 0.931 0.893 0.843 0.031 0.937
VSCode 2D Prompt Swin 0.873 0.820 0.046 0.925 0.869 0.780 0.025 0.931 0.882 0.841 0.032 0.935
CGCOD Camoclass CLIP+PVTv2-b2 0.896 0.874 0.036 0.947 0.890 0.824 0.018 0.948 0.904 0.879 0.026 0.949
SFGNet Semantics CLIP+PVTv2-b2 0.911 0.875 0.035 0.944 0.887 0.823 0.019 0.940 0.907 0.879 0.025 0.952
SFGNet Semantics CLIP+PVTv2-b4 0.913 0.877 0.033 0.947 0.889 0.823 0.018 0.948 0.909 0.881 0.023 0.955
SFGNet(Ours) Semantics CLIP+PVTv2-b5 0.915 0.880 0.032 0.951 0.890 0.824 0.016 0.949 0.912 0.884 0.021 0.957
@@ 3.2 Comparison with State-of-the-art Methods
We compared SFGNet with several recent and representative methods, including PFNet [6], ZoomNet [18], FSPNet [19], FEDER [5], HitNet [20], FSEL [21], CamoFormer [8], FHCNet [22], VSCode [9], and CGCOD [10]. To ensure fair comparison, all prediction maps were either provided by the original authors or generated using the released codes. Table 1 reports the quantitative comparison between SFGNet and other models. Specifically, SFGNet outperforms the compared methods on CAMO, and NC4K datasets, while achieving comparable $S_{m}$ and $F_{\beta}^{\omega}$ performance with CGCOD on COD10K. Qualitative comparisons are illustrated in Figure 3, which demonstrate that SFGNet performs better than state-of-the-art approaches in most cases, and is comparable in a few challenging scenarios. In addition, we evaluated different PVT backbones and found that pvt-v2-b5 achieves the best performance.
@@ 3.3.1 Analysis of different semantic prompts
”A photo of camouflaged <class>.” 0.887 0.821 0.018 0.943
@@
”Camouflaged <class> blending
into the background.” 0.888 0.821 0.018 0.946
@@
surrounding environment.” (Ours) 0.890 0.824 0.016 0.949

### v2

@@ Abstract
Camouflaged object detection (COD) aims to segment objects that blend into their surroundings. However, most existing studies overlook the semantic differences among textual prompts of different targets as well as fine-grained frequency features. In this work, we propose a novel Semantic and Frequency Guided Network (SFGNet), which incorporates semantic prompts and frequency-domain features to capture camouflaged objects and improve boundary perception. We further design Multi-Band Fourier Module(MBFM) to enhance the ability of the network in handling complex backgrounds and blurred boundaries. In addition, we design an Interactive Structure Enhancement Block (ISEB) to ensure structural integrity and boundary details in the predictions. Extensive experiments conducted on three COD benchmark datasets demonstrate that our method significantly outperforms state-of-the-art approaches. The code will be released once the paper is accepted.
@@ 3.1.1 Datasets
PFNet [16] None ResNet 0.782 0.695 0.085 0.855 0.800 0.660 0.040 0.877 0.829 0.745 0.053 0.887
ZoomNet [17] None ResNet 0.820 0.752 0.066 0.877 0.838 0.729 0.029 0.888 0.853 0.784 0.043 0.896
FSPNet [18] None ViT 0.851 0.802 0.056 0.905 0.850 0.755 0.028 0.912 0.879 0.816 0.035 0.914
FEDER [19] None Res2Net 0.836 0.807 0.066 0.897 0.844 0.748 0.029 0.911 0.862 0.824 0.042 0.913
HitNet [20] None PVT 0.849 0.809 0.055 0.906 0.871 0.806 0.023 0.935 0.875 0.834 0.037 0.926
FSEL [21] None PVT 0.885 0.857 0.040 0.942 0.877 0.799 0.021 0.937 0.892 0.852 0.030 0.941
CamoFormer [6] None PVT 0.872 0.831 0.046 0.929 0.869 0.786 0.023 0.932 0.892 0.847 0.030 0.939
FHCNet [22] None Swin+Res2Net 0.885 0.842 0.039 0.934 0.869 0.778 0.023 0.931 0.893 0.843 0.031 0.937
VSCode [7] 2D Prompt Swin 0.873 0.820 0.046 0.925 0.869 0.780 0.025 0.931 0.882 0.841 0.032 0.935
CGCOD [8] Camoclass CLIP+PVT 0.896 0.874 0.036 0.947 0.890 0.824 0.018 0.948 0.904 0.879 0.026 0.949
SFGNet Semantics CLIP+PVTv2-b2 0.904 0.878 0.034 0.949 0.890 0.826 0.019 0.947 0.907 0.879 0.025 0.952
SFGNet Semantics CLIP+PVTv2-b4 0.906 0.879 0.033 0.949 0.890 0.827 0.018 0.948 0.909 0.881 0.023 0.955
SFGNet(Ours) Semantics CLIP+PVTv2-b5 0.906 0.880 0.032 0.951 0.891 0.829 0.018 0.949 0.912 0.884 0.021 0.957
@@ 3.2 Comparison with State-of-the-art Methods
We compared SFGNet with several recent and representative methods, including PFNet [16], ZoomNet [17], FSPNet [18], FEDER [19], HitNet [20], FSEL [21], CamoFormer [6], FHCNet [22], VSCode [7], and CGCOD [8]. To ensure fair comparison, all prediction maps were either provided by the original authors or generated using the released codes. Table 1 reports the quantitative comparison between SFGNet and other models. Specifically, SFGNet outperforms the compared methods on CAMO, and NC4K datasets, while achieving comparable performance with CGCOD on COD10K. Qualitative comparisons are illustrated in Figure 3, which demonstrate that SFGNet performs better than state-of-the-art approaches in most cases, and is comparable in a few challenging scenarios. In addition, we evaluated different PVT backbones and found that pvt-v2-b5 achieves the best performance.
@@ 3.3.1 Analysis of different semantic prompts
”A photo of camouflaged <class>.” 0.887 0.824 0.019 0.943
@@
”Camouflaged object naturally blending
into the surrounding environment.” 0.888 0.824 0.019 0.946
@@
surrounding environment.” (Ours) 0.891 0.829 0.018 0.949